# Nonlinear Concept Erasure: a Density Matching Approach

### v1

@@ 5.3 Fairness Analysis
Fairness evaluation results, reported in Table 2, demonstrate that $\overline{\mathrm{L}}$ significantly reduces gender and racial bias in downstream classification tasks, achieving performance comparable to the strongest baselines at low rank (e.g., $r=10$ and $r=50$ with standard and cascaded training, respectively). The bias metrics consistently decrease as the number of preserved dimensions in the feature space is reduced, indicating an increasing level of bias mitigation.
Table 2(a) highlights a trade-off specific to the Bias in Bios dataset between classifier performance on the downstream task (occupation prediction) and classifier fairness. [^2] Such a trade-off has been demonstrated by Zaho & Gordon (2022) For FaRM and KRaM, achieving fairer classifiers comes at the cost of a substantial drop in downstream task accuracy. In contrast, the tunable rank in $\overline{\mathrm{L}}$ enables a more flexible compromise—for example, achieving significant gender bias reduction while incurring a smaller loss in downstream performance at $r=350$ or $r=250$ with Cascaded $\overline{\mathrm{L}}$. The significant reduction of the positive correlation between the TPR gender gap for an occupation and the gender imbalance in that occupation at $r=250$ is made clearly visible in Figure 4.
The results in Table 2(b) reveal a more nuanced picture: although $\overline{\mathrm{L}}$ outperforms the baselines in bias reduction at low rank (e.g. $r=15$), it does so at the expense of lower downstream accuracy. Depending on the application context, particularly when fairness is a primary concern, the ability to adjust the rank makes $\overline{\mathrm{L}}$ a favorable choice over less flexible methods.
@@
Figure 4: Average $\mathrm{TPR}\text{-Gap}_{\text{female},y}$ vs. female proportion in each occupation $y$ for different representations used train the classifiers. $P$ is the rank-250 projection calculated via Cascaded $\overline{\mathrm{L}}$. Each vertical corresponds to a specific occupation (e.g., nurse).
@@ 5.4 About MMD design choices
This work proposes an implementation of $\overline{\mathrm{L}}$ designed for broad applicability with minimal hyperparameter tuning. As detailed in Section 4.2, several choices regarding the evaluation of MMD were informed by prior studies and supported by empirical evidence. Specifically, employing a single kernel ($M=1$) yields results of comparable magnitude to those reported in this article, while increasing $M$ beyond 2 kernels yields only marginal improvements (see Supplementary Material F). We set $M=5$ in all experiments as a trade-off between performance and computational cost.
@@ 5.5 About the freedom of choice for the rank
There is no universally optimal rank $r$ for concept erasure with $\overline{\mathrm{L}}$. The ability to adjust the rank is a key strength of the method, allowing for context-specific trade-offs between erasure effectiveness and post-erasure representation utility. In practice, several strategies can guide the selection of an appropriate rank.
@@ 6 Scope and Limitations
Our method specifically targets the erasure of discrete concepts. It could be naturally extended to the erasure of multiple discrete concepts by replacing the erasure loss (Eq. 6) with a sum of individual erasure losses, one per concept to be erased, in Eq. 7. With $\overline{\mathrm{L}}$, we do not address the erasure of continuous concepts, as the density matching objective in Eq. 6 requires partitioning samples according to discrete concept classes. For continuous concept erasure, approaches such as FaRM [3] or KRaM [4] are more suitable.
@@ 7 Conclusion and future work
We introduced $\overline{\mathrm{L}}$, a simple yet effective method for nonlinear concept erasure which uses orthogonal projections to neutralize concept-specific information. By optimizing a density matching objective, $\overline{\mathrm{L}}$ ensures indistinguishability of class-conditional distributions in the the embedding space after erasure. Evaluated on classic NLP benchmarks, $\overline{\mathrm{L}}$ is competitive and even outperforms state-of-the-art methods in nonlinear erasure of demographic attributes while better preserving the post-erasure utility of the representations. A notable strength of $\overline{\mathrm{L}}$ is its ability to control the trade-off between concept erasure and information preservation by adjusting the rank of the learned projection. Empirical results also demonstrate that $\overline{\mathrm{L}}$ improves fairness in downstream classifiers with a moderate impact on task performance.
@@ Appendix C Training Details
Optimization and Hyperparameters for $\overline{\mathrm{\textbf{L}}}$
@@
  • $\overline{\mathrm{L}}$: $\gamma$ is scaled as $\frac{1}{r^{2}}$, using $\frac{200}{r^{2}}$ for GloVe embeddings and $\frac{100}{r^{2}}$ for Bias in Bios and Dial.
@@ D.1 GloVe
$a_{z}^{\mathrm{bin}}$downarrow (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$downarrow (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$ Rank
@@
$\overline{\mathrm{L}}$ (standard)
||||  $r=275$ 88.8 $\pm$ 0.5 7.1 70.0 $\pm$ 0.4 0.88 0.72 275
||||  $r=250$ 85.5 $\pm$ 0.3 8.7 65.5 $\pm$ 0.6 0.86 0.69 250
||||  $r=225$ 82.1 $\pm$ 0.5 10.2 60.9 $\pm$ 0.6 0.85 0.66 225
||||  $r=200$ 77.5 $\pm$ 0.5 12.0 56.6 $\pm$ 0.4 0.83 0.65 200
||||  $r=175$ 74.0 $\pm$ 0.9 13.6 53.3 $\pm$ 0.6 0.82 0.62 175
||||  $r=150$ 70.3 $\pm$ 0.5 15.3 49.7 $\pm$ 1.2 0.81 0.58 150
||||  $r=125$ 65.6 $\pm$ 0.5 18.3 46.2 $\pm$ 0.5 0.80 0.54 125
||||  $r=100$ 61.4 $\pm$ 0.7 21.7 42.7 $\pm$ 0.7 0.78 0.48 100
||||  $r=75$ 57.3 $\pm$ 0.6 26.2 39.7 $\pm$ 0.4 0.75 0.45 75
||||  $r=50$ 54.1 $\pm$ 0.8 28.3 37.0 $\pm$ 0.7 0.72 0.34 50
||||  $r=25$ 51.1 $\pm$ 0.8 19.3 35.4 $\pm$ 0.8 0.68 0.24 25
$\overline{\mathrm{L}}$ (cascaded)
||||  $r=275$ 87.3 $\pm$ 0.4 8.0 68.1 $\pm$ 0.7 0.87 0.73 275
||||  $r=250$ 80.6 $\pm$ 0.2 11.6 60.2 $\pm$ 0.6 0.84 0.68 250
||||  $r=225$ 71.4 $\pm$ 0.8 15.7 51.0 $\pm$ 0.5 0.82 0.65 225
||||  $r=200$ 65.2 $\pm$ 0.7 19.8 45.4 $\pm$ 0.7 0.81 0.64 200
||||  $r=175$ 58.9 $\pm$ 0.4 23.8 41.6 $\pm$ 0.3 0.79 0.63 175
||||  $r=150$ 55.9 $\pm$ 1.1 27.1 38.1 $\pm$ 0.6 0.78 0.64 150
||||  $r=125$ 51.8 $\pm$ 0.9 30.4 36.3 $\pm$ 0.5 0.76 0.64 125
||||  $r=100$ 51.0 $\pm$ 0.4 32.4 35.4 $\pm$ 0.7 0.74 0.64 100
||||  $r=75$ 50.7 $\pm$ 0.7 34.1 34.2 $\pm$ 0.7 0.73 0.60 75
||||  $r=50$ 50.4 $\pm$ 0.6 34.0 34.3 $\pm$ 1.0 0.70 0.60 50
||||  $r=25$ 50.1 $\pm$ 0.9 25.9 34.2 $\pm$ 0.8 0.66 0.49 25
Figure 5: t-SNE visualization of GloVe embeddings before erasure (left) and after erasure via $\overline{\mathrm{L}}$ with a rank-150 orthogonal projection (right). Male-biased, female-biased, and neutral words are displayed in different colors.
@@ D.2 Bias in Bios
$a_{z}$downarrow (%) MDL $a_{y}$uparrow (%) Rank
@@
$\overline{\mathrm{L}}$
||||  $r=750$ 98.4 $\pm$ 0.1 31.0 79.6 $\pm$ 0.2 750
||||  $r=700$ 97.0 $\pm$ 0.9 55.7 79.5 $\pm$ 0.2 700
||||  $r=650$ 96.7 $\pm$ 0.4 77.4 79.4 $\pm$ 0.2 650
||||  $r=600$ 95.6 $\pm$ 0.2 90.9 79.1 $\pm$ 0.3 600
||||  $r=550$ 94.3 $\pm$ 1.0 97.5 79.0 $\pm$ 0.2 550
||||  $r=500$ 93.4 $\pm$ 0.8 112.0 78.9 $\pm$ 0.1 500
||||  $r=450$ 92.8 $\pm$ 0.6 110.7 78.6 $\pm$ 0.4 450
||||  $r=400$ 91.6 $\pm$ 0.8 119.1 78.4 $\pm$ 0.3 400
||||  $r=350$ 90.1 $\pm$ 0.8 126.4 78.2 $\pm$ 0.1 350
||||  $r=300$ 88.9 $\pm$ 0.3 128.3 77.9 $\pm$ 0.1 300
||||  $r=250$ 88.2 $\pm$ 0.6 133.8 77.4 $\pm$ 0.2 250
||||  $r=200$ 85.7 $\pm$ 0.4 142.3 76.8 $\pm$ 0.1 200
||||  $r=150$ 82.1 $\pm$ 0.4 158.7 75.5 $\pm$ 0.2 150
||||  $r=100$ 77.1 $\pm$ 0.7 172.0 73.2 $\pm$ 0.2 100
||||  $r=50$ 69.0 $\pm$ 0.7 186.2 67.9 $\pm$ 0.2 50
||||  $r=50$ 58.8 $\pm$ 0.7 193.4 53.4 $\pm$ 0.1 10
$\overline{\mathrm{L}}$ (cascaded)
||||  $r=750$ 97.1 $\pm$ 0.2 48.7 79.1 $\pm$ 0.2 750
||||  $r=700$ 93.7 $\pm$ 0.4 95.7 78.5 $\pm$ 0.2 700
||||  $r=650$ 90.6 $\pm$ 0.4 126.7 77.6 $\pm$ 0.1 650
||||  $r=600$ 86.3 $\pm$ 1.2 147.3 76.7 $\pm$ 0.1 600
||||  $r=550$ 82.8 $\pm$ 1.7 160.7 76.2 $\pm$ 0.2 550
||||  $r=500$ 79.7 $\pm$ 1.5 171.5 75.5 $\pm$ 0.1 500
||||  $r=450$ 76.2 $\pm$ 1.8 176.0 74.7 $\pm$ 0.2 450
||||  $r=400$ 72.6 $\pm$ 2.6 182.2 73.7 $\pm$ 0.3 400
||||  $r=350$ 67.4 $\pm$ 2.0 185.5 72.7 $\pm$ 0.3 350
||||  $r=300$ 63.1 $\pm$ 1.4 189.8 71.7 $\pm$ 0.4 300
||||  $r=250$ 59.7 $\pm$ 0.6 190.8 70.1 $\pm$ 0.4 250
||||  $r=200$ 57.3 $\pm$ 0.9 191.2 68.1 $\pm$ 0.6 200
||||  $r=150$ 55.9 $\pm$ 0.3 192.2 65.5 $\pm$ 0.5 150
||||  $r=100$ 55.1 $\pm$ 0.3 193.1 61.5 $\pm$ 0.6 100
||||  $r=50$ 54.1 $\pm$ 0.2 193.5 54.1 $\pm$ 0.6 50
Figure 6: t-SNE visualization of Bias in Bios embeddings before erasure (left) and after erasure via cascaded $\overline{\mathrm{L}}$ with a rank-250 orthogonal projection (right). Female-biased and male-biased text representations are displayed in different colors.
@@ D.3 DIAL
$a_{z}$downarrow MDL $\uparrow$ $a_{y}$uparrow DP $\downarrow$ rank
@@
$\overline{\mathrm{L}}$
||||  $r=275$ 86.3 $\pm$ 0.2 176.0 75.5 $\pm$ 0.2 0.19 275
||||  $r=250$ 86.0 $\pm$ 0.3 183.1 75.3 $\pm$ 0.2 0.21 250
||||  $r=225$ 85.4 $\pm$ 0.2 193.1 75.3 $\pm$ 0.1 0.21 225
||||  $r=200$ 84.8 $\pm$ 0.2 200.6 75.2 $\pm$ 0.2 0.18 200
||||  $r=175$ 83.8 $\pm$ 0.4 210.7 74.9 $\pm$ 0.3 0.17 175
||||  $r=150$ 82.8 $\pm$ 0.2 223.4 74.9 $\pm$ 0.1 0.17 150
||||  $r=125$ 80.9 $\pm$ 0.3 238.6 74.1 $\pm$ 0.2 0.15 125
||||  $r=100$ 78.9 $\pm$ 0.5 248.3 74.4 $\pm$ 0.1 0.18 100
||||  $r=75$ 75.8 $\pm$ 0.3 269.3 73.8 $\pm$ 0.2 0.13 75
||||  $r=50$ 70.2 $\pm$ 0.7 290.7 71.6 $\pm$ 0.2 0.13 50
||||  $r=40$ 67.8 $\pm$ 0.2 298.3 70.4 $\pm$ 0.2 0.09 40
||||  $r=25$ 61.1 $\pm$ 0.4 307.9 68.4 $\pm$ 0.1 0.04 25
||||  $r=15$ 56.8 $\pm$ 0.2 311.3 67.4 $\pm$ 0.1 0.02 15
$\overline{\mathrm{L}}$ cascaded
||||  $r=275$ 85.5 $\pm$ 0.2 189.6 75.1 $\pm$ 0.1 0.17 275
||||  $r=250$ 84.5 $\pm$ 0.3 204.1 74.8 $\pm$ 0.2 0.16 250
||||  $r=225$ 83.8 $\pm$ 0.3 213.7 74.8 $\pm$ 0.1 0.15 225
||||  $r=200$ 83.1 $\pm$ 0.2 223.9 74.8 $\pm$ 0.1 0.15 200
||||  $r=175$ 81.3 $\pm$ 0.2 236.0 74.6 $\pm$ 0.3 0.13 175
||||  $r=150$ 80.4 $\pm$ 0.4 245.0 74.7 $\pm$ 0.2 0.13 150
||||  $r=125$ 78.5 $\pm$ 0.2 257.0 74.1 $\pm$ 0.2 0.14 125
||||  $r=100$ 77.4 $\pm$ 0.3 264.1 74.4 $\pm$ 0.2 0.14 100
||||  $r=75$ 73.3 $\pm$ 0.4 281.5 73.5 $\pm$ 0.2 0.12 75
||||  $r=50$ 67.6 $\pm$ 0.3 297.6 71.8 $\pm$ 0.2 0.08 50
||||  $r=40$ 64.5 $\pm$ 0.4 303.3 70.8 $\pm$ 0.4 0.07 40
||||  $r=25$ 62.4 $\pm$ 0.3 303.0 72.8 $\pm$ 0.2 0.11 25
||||  $r=15$ 57.2 $\pm$ 0.3 309.0 70.2 $\pm$ 0.1 0.07 15
Figure 7: t-SNE visualization of DIAL embeddings before erasure (left) and after erasure via cascaded $\overline{\mathrm{L}}$ with a rank-50 orthogonal projection (right). Representations of tweets written in African-American English and Standard American English are displayed in different colors.

### v2

@@ 5.3 Fairness Analysis
Fairness evaluation results, reported in Table 2, demonstrate that $\overline{\mathrm{L}}$EOPARD significantly reduces gender and racial bias in downstream classification tasks, achieving performance comparable to the strongest baselines at low rank (e.g., $r=10$ and $r=50$ with standard and cascaded training, respectively). The bias metrics consistently decrease as the number of preserved dimensions in the feature space is reduced, indicating an increasing level of bias mitigation.
Table 2(a) highlights a trade-off specific to the Bias in Bios dataset between classifier performance on the downstream task (occupation prediction) and classifier fairness. [^2] Such a trade-off has been demonstrated by Zaho & Gordon (2022) For FaRM and KRaM, achieving fairer classifiers comes at the cost of a substantial drop in downstream task accuracy. In contrast, the tunable rank in $\overline{\mathrm{L}}$EOPARD enables a more flexible compromise—for example, achieving significant gender bias reduction while incurring a smaller loss in downstream performance at $r=350$ or $r=250$ with Cascaded $\overline{\mathrm{L}}$EOPARD. The significant reduction of the positive correlation between the TPR gender gap for an occupation and the gender imbalance in that occupation at $r=250$ is made clearly visible in Figure 4.
The results in Table 2(b) reveal a more nuanced picture: although $\overline{\mathrm{L}}$EOPARD outperforms the baselines in bias reduction at low rank (e.g. $r=15$), it does so at the expense of lower downstream accuracy. Depending on the application context, particularly when fairness is a primary concern, the ability to adjust the rank makes $\overline{\mathrm{L}}$EOPARD a favorable choice over less flexible methods.
@@
Figure 4: Average $\mathrm{TPR}\text{-Gap}_{\text{female},y}$ vs. female proportion in each occupation $y$ for different representations used train the classifiers. $P$ is the rank-250 projection calculated via Cascaded $\overline{\mathrm{L}}$EOPARD. Each vertical corresponds to a specific occupation (e.g., nurse).
@@ 5.4 About MMD design choices
This work proposes an implementation of $\overline{\mathrm{L}}$EOPARD designed for broad applicability with minimal hyperparameter tuning. As detailed in Section 4.2, several choices regarding the evaluation of MMD were informed by prior studies and supported by empirical evidence. Specifically, employing a single kernel ($M=1$) yields results of comparable magnitude to those reported in this article, while increasing $M$ beyond 2 kernels yields only marginal improvements (see Supplementary Material F). We set $M=5$ in all experiments as a trade-off between performance and computational cost.
@@ 5.5 About the freedom of choice for the rank
There is no universally optimal rank $r$ for concept erasure with $\overline{\mathrm{L}}$EOPARD. The ability to adjust the rank is a key strength of the method, allowing for context-specific trade-offs between erasure effectiveness and post-erasure representation utility. In practice, several strategies can guide the selection of an appropriate rank.
@@ 6 Scope and Limitations
Our method specifically targets the erasure of discrete concepts. It could be naturally extended to the erasure of multiple discrete concepts by replacing the erasure loss (Eq. 6) with a sum of individual erasure losses, one per concept to be erased, in Eq. 7. With $\overline{\mathrm{L}}$EOPARD, we do not address the erasure of continuous concepts, as the density matching objective in Eq. 6 requires partitioning samples according to discrete concept classes. For continuous concept erasure, approaches such as FaRM [3] or KRaM [4] are more suitable.
@@ 7 Conclusion and future work
We introduced $\overline{\mathrm{L}}$EOPARD, a simple yet effective method for nonlinear concept erasure which uses orthogonal projections to neutralize concept-specific information. By optimizing a density matching objective, $\overline{\mathrm{L}}$EOPARD ensures indistinguishability of class-conditional distributions in the the embedding space after erasure. Evaluated on classic NLP benchmarks, $\overline{\mathrm{L}}$EOPARD is competitive and even outperforms state-of-the-art methods in nonlinear erasure of demographic attributes while better preserving the post-erasure utility of the representations. A notable strength of $\overline{\mathrm{L}}$EOPARD is its ability to control the trade-off between concept erasure and information preservation by adjusting the rank of the learned projection. Empirical results also demonstrate that $\overline{\mathrm{L}}$EOPARD improves fairness in downstream classifiers with a moderate impact on task performance.
@@ Appendix C Training Details
Optimization and Hyperparameters for $\overline{\mathrm{\textbf{L}}}$EOPARD
@@
  • $\overline{\mathrm{L}}$EOPARD: $\gamma$ is scaled as $\frac{1}{r^{2}}$, using $\frac{200}{r^{2}}$ for GloVe embeddings and $\frac{100}{r^{2}}$ for Bias in Bios and Dial.
@@ D.1 GloVe
$a_{z}^{\mathrm{bin}}$ $\downarrow$ (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$ $\downarrow$ (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$ Rank
@@
$\overline{\mathrm{L}}$EOPARD (standard)
$|$  $r=275$ 88.8 $\pm$ 0.5 7.1 70.0 $\pm$ 0.4 0.88 0.72 275
$|$  $r=250$ 85.5 $\pm$ 0.3 8.7 65.5 $\pm$ 0.6 0.86 0.69 250
$|$  $r=225$ 82.1 $\pm$ 0.5 10.2 60.9 $\pm$ 0.6 0.85 0.66 225
$|$  $r=200$ 77.5 $\pm$ 0.5 12.0 56.6 $\pm$ 0.4 0.83 0.65 200
$|$  $r=175$ 74.0 $\pm$ 0.9 13.6 53.3 $\pm$ 0.6 0.82 0.62 175
$|$  $r=150$ 70.3 $\pm$ 0.5 15.3 49.7 $\pm$ 1.2 0.81 0.58 150
$|$  $r=125$ 65.6 $\pm$ 0.5 18.3 46.2 $\pm$ 0.5 0.80 0.54 125
$|$  $r=100$ 61.4 $\pm$ 0.7 21.7 42.7 $\pm$ 0.7 0.78 0.48 100
$|$  $r=75$ 57.3 $\pm$ 0.6 26.2 39.7 $\pm$ 0.4 0.75 0.45 75
$|$  $r=50$ 54.1 $\pm$ 0.8 28.3 37.0 $\pm$ 0.7 0.72 0.34 50
$|$  $r=25$ 51.1 $\pm$ 0.8 19.3 35.4 $\pm$ 0.8 0.68 0.24 25
$\overline{\mathrm{L}}$EOPARD (cascaded)
$|$  $r=275$ 87.3 $\pm$ 0.4 8.0 68.1 $\pm$ 0.7 0.87 0.73 275
$|$  $r=250$ 80.6 $\pm$ 0.2 11.6 60.2 $\pm$ 0.6 0.84 0.68 250
$|$  $r=225$ 71.4 $\pm$ 0.8 15.7 51.0 $\pm$ 0.5 0.82 0.65 225
$|$  $r=200$ 65.2 $\pm$ 0.7 19.8 45.4 $\pm$ 0.7 0.81 0.64 200
$|$  $r=175$ 58.9 $\pm$ 0.4 23.8 41.6 $\pm$ 0.3 0.79 0.63 175
$|$  $r=150$ 55.9 $\pm$ 1.1 27.1 38.1 $\pm$ 0.6 0.78 0.64 150
$|$  $r=125$ 51.8 $\pm$ 0.9 30.4 36.3 $\pm$ 0.5 0.76 0.64 125
$|$  $r=100$ 51.0 $\pm$ 0.4 32.4 35.4 $\pm$ 0.7 0.74 0.64 100
$|$  $r=75$ 50.7 $\pm$ 0.7 34.1 34.2 $\pm$ 0.7 0.73 0.60 75
$|$  $r=50$ 50.4 $\pm$ 0.6 34.0 34.3 $\pm$ 1.0 0.70 0.60 50
$|$  $r=25$ 50.1 $\pm$ 0.9 25.9 34.2 $\pm$ 0.8 0.66 0.49 25
Figure 5: t-SNE visualization of GloVe embeddings before erasure (left) and after erasure via $\overline{\mathrm{L}}$EOPARD with a rank-150 orthogonal projection (right). Male-biased, female-biased, and neutral words are displayed in different colors.
@@ D.2 Bias in Bios
$a_{z}$ $\downarrow$ (%) MDL $a_{y}$ $\uparrow$ (%) Rank
@@
$\overline{\mathrm{L}}$EOPARD
$|$  $r=750$ 98.4 $\pm$ 0.1 31.0 79.6 $\pm$ 0.2 750
$|$  $r=700$ 97.0 $\pm$ 0.9 55.7 79.5 $\pm$ 0.2 700
$|$  $r=650$ 96.7 $\pm$ 0.4 77.4 79.4 $\pm$ 0.2 650
$|$  $r=600$ 95.6 $\pm$ 0.2 90.9 79.1 $\pm$ 0.3 600
$|$  $r=550$ 94.3 $\pm$ 1.0 97.5 79.0 $\pm$ 0.2 550
$|$  $r=500$ 93.4 $\pm$ 0.8 112.0 78.9 $\pm$ 0.1 500
$|$  $r=450$ 92.8 $\pm$ 0.6 110.7 78.6 $\pm$ 0.4 450
$|$  $r=400$ 91.6 $\pm$ 0.8 119.1 78.4 $\pm$ 0.3 400
$|$  $r=350$ 90.1 $\pm$ 0.8 126.4 78.2 $\pm$ 0.1 350
$|$  $r=300$ 88.9 $\pm$ 0.3 128.3 77.9 $\pm$ 0.1 300
$|$  $r=250$ 88.2 $\pm$ 0.6 133.8 77.4 $\pm$ 0.2 250
$|$  $r=200$ 85.7 $\pm$ 0.4 142.3 76.8 $\pm$ 0.1 200
$|$  $r=150$ 82.1 $\pm$ 0.4 158.7 75.5 $\pm$ 0.2 150
$|$  $r=100$ 77.1 $\pm$ 0.7 172.0 73.2 $\pm$ 0.2 100
$|$  $r=50$ 69.0 $\pm$ 0.7 186.2 67.9 $\pm$ 0.2 50
$|$  $r=10$ 58.8 $\pm$ 0.7 193.4 53.4 $\pm$ 0.1 10
$\overline{\mathrm{L}}$EOPARD (cascaded)
$|$  $r=750$ 97.1 $\pm$ 0.2 48.7 79.1 $\pm$ 0.2 750
$|$  $r=700$ 93.7 $\pm$ 0.4 95.7 78.5 $\pm$ 0.2 700
$|$  $r=650$ 90.6 $\pm$ 0.4 126.7 77.6 $\pm$ 0.1 650
$|$  $r=600$ 86.3 $\pm$ 1.2 147.3 76.7 $\pm$ 0.1 600
$|$  $r=550$ 82.8 $\pm$ 1.7 160.7 76.2 $\pm$ 0.2 550
$|$  $r=500$ 79.7 $\pm$ 1.5 171.5 75.5 $\pm$ 0.1 500
$|$  $r=450$ 76.2 $\pm$ 1.8 176.0 74.7 $\pm$ 0.2 450
$|$  $r=400$ 72.6 $\pm$ 2.6 182.2 73.7 $\pm$ 0.3 400
$|$  $r=350$ 67.4 $\pm$ 2.0 185.5 72.7 $\pm$ 0.3 350
$|$  $r=300$ 63.1 $\pm$ 1.4 189.8 71.7 $\pm$ 0.4 300
$|$  $r=250$ 59.7 $\pm$ 0.6 190.8 70.1 $\pm$ 0.4 250
$|$  $r=200$ 57.3 $\pm$ 0.9 191.2 68.1 $\pm$ 0.6 200
$|$  $r=150$ 55.9 $\pm$ 0.3 192.2 65.5 $\pm$ 0.5 150
$|$  $r=100$ 55.1 $\pm$ 0.3 193.1 61.5 $\pm$ 0.6 100
$|$  $r=50$ 54.1 $\pm$ 0.2 193.5 54.1 $\pm$ 0.6 50
Figure 6: t-SNE visualization of Bias in Bios embeddings before erasure (left) and after erasure via cascaded $\overline{\mathrm{L}}$EOPARD with a rank-250 orthogonal projection (right). Female-biased and male-biased text representations are displayed in different colors.
@@ D.3 DIAL
$a_{z}$ $\downarrow$ MDL $\uparrow$ $a_{y}$ $\uparrow$ DP $\downarrow$ rank
@@
$\overline{\mathrm{L}}$EOPARD
$|$  $r=275$ 86.3 $\pm$ 0.2 176.0 75.5 $\pm$ 0.2 0.19 275
$|$  $r=250$ 86.0 $\pm$ 0.3 183.1 75.3 $\pm$ 0.2 0.21 250
$|$  $r=225$ 85.4 $\pm$ 0.2 193.1 75.3 $\pm$ 0.1 0.21 225
$|$  $r=200$ 84.8 $\pm$ 0.2 200.6 75.2 $\pm$ 0.2 0.18 200
$|$  $r=175$ 83.8 $\pm$ 0.4 210.7 74.9 $\pm$ 0.3 0.17 175
$|$  $r=150$ 82.8 $\pm$ 0.2 223.4 74.9 $\pm$ 0.1 0.17 150
$|$  $r=125$ 80.9 $\pm$ 0.3 238.6 74.1 $\pm$ 0.2 0.15 125
$|$  $r=100$ 78.9 $\pm$ 0.5 248.3 74.4 $\pm$ 0.1 0.18 100
$|$  $r=75$ 75.8 $\pm$ 0.3 269.3 73.8 $\pm$ 0.2 0.13 75
$|$  $r=50$ 70.2 $\pm$ 0.7 290.7 71.6 $\pm$ 0.2 0.13 50
$|$  $r=40$ 67.8 $\pm$ 0.2 298.3 70.4 $\pm$ 0.2 0.09 40
$|$  $r=25$ 61.1 $\pm$ 0.4 307.9 68.4 $\pm$ 0.1 0.04 25
$|$  $r=15$ 56.8 $\pm$ 0.2 311.3 67.4 $\pm$ 0.1 0.02 15
$\overline{\mathrm{L}}$EOPARD cascaded
$|$  $r=275$ 85.5 $\pm$ 0.2 189.6 75.1 $\pm$ 0.1 0.17 275
$|$  $r=250$ 84.5 $\pm$ 0.3 204.1 74.8 $\pm$ 0.2 0.16 250
$|$  $r=225$ 83.8 $\pm$ 0.3 213.7 74.8 $\pm$ 0.1 0.15 225
$|$  $r=200$ 83.1 $\pm$ 0.2 223.9 74.8 $\pm$ 0.1 0.15 200
$|$  $r=175$ 81.3 $\pm$ 0.2 236.0 74.6 $\pm$ 0.3 0.13 175
$|$  $r=150$ 80.4 $\pm$ 0.4 245.0 74.7 $\pm$ 0.2 0.13 150
$|$  $r=125$ 78.5 $\pm$ 0.2 257.0 74.1 $\pm$ 0.2 0.14 125
$|$  $r=100$ 77.4 $\pm$ 0.3 264.1 74.4 $\pm$ 0.2 0.14 100
$|$  $r=75$ 73.3 $\pm$ 0.4 281.5 73.5 $\pm$ 0.2 0.12 75
$|$  $r=50$ 67.6 $\pm$ 0.3 297.6 71.8 $\pm$ 0.2 0.08 50
$|$  $r=40$ 64.5 $\pm$ 0.4 303.3 70.8 $\pm$ 0.4 0.07 40
$|$  $r=25$ 62.4 $\pm$ 0.3 303.0 72.8 $\pm$ 0.2 0.11 25
$|$  $r=15$ 57.2 $\pm$ 0.3 309.0 70.2 $\pm$ 0.1 0.07 15
Figure 7: t-SNE visualization of DIAL embeddings before erasure (left) and after erasure via cascaded $\overline{\mathrm{L}}$EOPARD with a rank-50 orthogonal projection (right). Representations of tweets written in African-American English and Standard American English are displayed in different colors.